OpenStudy (anonymous):

Solve the equation: x2 - 10x + 25 = 16. Enter the smallest zero value. Use a number only. is it 4.5?

OpenStudy (anonymous):

No. 4.5^2-10(4.5)+25=0.25, not 16

OpenStudy (anonymous):

that how the question was asked

OpenStudy (radar):

Just what is "the smallest zero value" You sure it didn't say "the smallest nonzero value"

OpenStudy (anonymous):

nope i copyed and pasted the question

OpenStudy (radar):

Just solve the equation. \[x ^{2}-10x+25=16\]\[x ^{2}-10x+9=0\]\[(x-9)(x-1)=0\]\[x=9\]\[x=1\]

OpenStudy (anonymous):

ok

OpenStudy (radar):

They both work, i guess the x=1 is closest to zero, if that is really what is wanted???

OpenStudy (anonymous):

yea i didnt type the question i copied and pasted it

OpenStudy (radar):

The question is kind of stooooopid lol

OpenStudy (anonymous):

lol